OpenStudy (anonymous):

Please enter a differential equation has a solution y = 3sin (4x + v). You do not need any boundary conditions

OpenStudy (anonymous):

\[\frac{\partial^2 y}{\partial x^2} = 16 \frac{\partial^2 y}{\partial v^2}\] works just fine, though it wouldn't specify the 3 without initial or boundary conditions...